# Learning to Learn for Few-shot Continual Active Learning

### v3

@@ 21             Stop training
   w/. Ring buffer (5 per class) 66.9
@@
   w/. Reservoir (default) 66.5
@@
FULL 64.1 $\pm$ 1.9
Outer-loop optimization
Outer-loop algorithm optimizes initial parameter set $\theta$ to a setting that $f$ can effectively adapt to $\mathcal{D}_{1:t}$ via a few gradient updates. Due to memory constraints, we only have a few amount of samples from the past. The model reads all examples from memory $\mathcal{M}$. Then, the outer-loop objective is to have $\tilde{\theta}=\theta_{\mathrm{e}}\cup\tilde{\theta}_{\mathrm{clf}}$ generalize well on $\mathcal{D}_{1:t}$ using $\mathcal{M}$, shown in Eqn. 6,
To reduce the complexity of the second-order computation in outer loop, we use first-order approximation, namely FOMAML. Then, we use Adam [27] for outer-loop optimization as
V-C Testing
where $S\subseteq\mathcal{M}$. And the optimization process is,
VI Experiments
VI-A Datasets
We use the text classification benchmark datasets from [5], including AGNews (news classification; 4 classes), Yelp (sentiment analysis; 5 classes), Amazon (sentiment analysis; 5 classes), DBpedia (Wikipedia article classification; 14 classes) and Yahoo (questions and answers categorization; 10 classes). This collection contains 5 tasks from 4 different domains, which covers class- and domain-incremental learning in task sequence. We randomly sample 5 labeled instances per class, 10,000 unlabeled instances, and 7,600 test examples from each of the datasets. Following prior studies [7, 8], we concatenate training sets in 4 different orderings as shown in Table II.
TABLE II: Input Datasets Orders
Index Dataset Orders
VI-B Implementation details
VI-C Baselines
Baseline model: (a) MAML-SEQ: online FOMAML algorithm. (b) OML-ER [8]: OML with 5% episodic experience replay rate [^3] The default 1% rate shows bad performance. (c) C-MAML [^4] Being modified for NLP tasks [4]: OML with Meta & CL objective alignment + Reservoir Sampling. (d) Meta-CAL (ours.): OML with Meta & CL objective alignment + consistency regularization + reservoir sampling. (e) FULL: supervised C-MAML, trained on full datasets. Memory sample selection: (f) Prototype [11]: select representative samples that are closest to dynamically updated prototypes in representation space. (g) Ring Buffer [30]: use a ‘First-In, First-Out’ scheme to update buffer. (h) Reservoir Sampling [3]: randomly select data with an equal selection probability.
VI-D Evaluation Metrics
Based on the prior work [31], we use three evaluation metrics, i.e., accuracy, backward transfer and forward transfer. Let $R_{k,i}$ be the macro-averaged accuracy evaluated on the test set of the $i$ task after sequentially learning $k$ tasks:
Accuracy
$\mathrm{ACC}_{k}=\frac{1}{k}\sum^{k}_{i=1}R_{k,i}$
Backward transfer
$\mathrm{BWT}_{k}=\frac{1}{k-1}\sum^{k-1}_{i=1}(R_{k,i}-R_{i,i})$
Forward transfer
$\mathrm{FWT}_{k}=\frac{1}{k-1}\sum^{k}_{i=2}(R_{i,i}-R_{0,i}$
TABLE III: Comparison of different AL strategies. We report the average accuracy of training sequence Yelp $\rightarrow$ AGNews $\rightarrow$ DBpedia $\rightarrow$ Amazon $\rightarrow$ Yahoo.
@@ VI-D Evaluation Metrics
VI-E Results
In Table I, we compare various models with four AL strategies. These strategies include Random (denoted as RAND), Representative via KMeans (denoted as REP), Diversity via KMeans (denoted as DIV), and Uncertainty via Least-confidence (denoted as UNC). The label budget $B_{A}$ = 2000 examples per task. Each record is the average of three best results from five runs. Save ratio is the ratio of saved samples to the total number of seen samples.
As for the memory sample selection schemes, Ring Buffer outperforms Reservoir Sampling by less than 1% in RAND and REP. Whereas, Reservoir Sampling exhibits smaller standard deviations, indicating its robustness to training set orders. Prototype Sampling has the worst performance. We perform further analysis in §VII-A.
Both Table I and Table III demonstrate that Random is the best AL strategy for our model. As shown in Table III, for both Diversity- and Representative-based AL strategies, KMeans is the most effective approach. The uncertainty-based methods comparatively perform poorly. CL models intend to revisit past samples that can bring generalization. However, annotating and replaying uncertain samples cannot bring enough generalization, which is not beneficial for knowledge consolidation. Hence, it harms model performance.
VII Further Analysis
(a) After Learning Task 1
(b) After Learning Task 2
(c) After Learning Task 3
(d) After Learning Task 4
(e) After Learning Task 5
(a) RAND
(b) REP (KMeans)
(c) DIV (KMeans)
(d) UNC (Least-confidence)
VII-A Memory Insight
AL strategies
As shown in Fig. 1a to 1d, memory data in RAND show a good data dispersion within the last task (i.e., intra-task generalization) and across multiple tasks (i.e., inter-task generalization), resulting in the best accuracy. In contrast, UNC in Fig. 1d shows subpar performance in inter-task generalization. The result validates UNC’s incapability to achieve generalization and preserve knowledge. Both REP and DIV exhibit lower accuracy compared to RAND, with DIV showing an obvious lower sparsity. Therefore, it is important to find an optimal balance between representativeness and diversity to ensure generalization.
Memory sample selection methods
In Fig. 1a, 1e and 1f, the choice of memory sample selection method also affects the model’s performance. Prototype Sampling selects representative memory samples, potentially resulting in relatively less sparsity across multiple tasks. In contrast, Reservoir and Ring Buffer sampling strategies introduce randomness into the memory sample selection process, achieving better inter-task generalization.
@@ VII-A Memory Insight
Different learning stages
Fig. 2 presents the T-SNE visualization of memory samples at different learning stages. As show in Figure 2a to 2b, when the proposed model learns a small number of tasks or encounters small label spaces, it focuses on ensuring intra-task generalization. However, as the number of seen tasks increases, the model shifts its focus towards ensuring inter-task generalization as shown in 2d to 2e. Consequently, we observe a phenomenon wherein the memory samples from the last tasks cluster together.
VII-B Stability & Plasticity
Next we test the performance of different AL strategies in terms of stability and plasiticity.
(a) BWT (Stability)
(b) FWT (Plasticity)
Stability
The negative BWT values indicate forgetting. As shown in Fig. 4a, RAND shows the least forgetting, which indicates the best stability on past tasks. However, all methods exhibit a large forgetting after sequentially learning Task 3. Task 3 (DBpedia, 14 classes) has a relatively large label space compared to Task 1 (Yelp, 5 classes) and Task 2 (AGNews, 4 classes). As shown in Fig. 3, the accuracy on DBpedia exceeds the accuracy on other tasks. Since MAML learns high-quality reusable features for fast adaptation [32], it tends to find shared representations that specifically benefit tasks with large label spaces.
Plasticity
A positive FWT demonstrates the model’s ability of zero-shot learning. Task 4 (Amazon) are from the same domain as Task 1 (Yelp). RAND, REP and DIV have a positive FWT value after learning Amazon. The result suggests a positive forward knowledge transfer to the task from the same domain. Moreover, it demonstrates the effectiveness of our model in domain-incremental learning when employed with these three AL strategies.
VII-C Effect of Augmentations
We conduct an ablation study to analyze textual augmentations and their key roles in Meta-CAL. We examine the effect of inner-loop augmentation and outer loop augmentation in Table IV. The query samples in the outer-loop constrain the model behaviour. We hypothesize that learning sufficient generalizability from query samples helps retain knowledge and prevent catastrophic forgetting. The results validate this hypothesis, as we observe a significant gain by using outer-loop augmentation to enhance generalization. It improves accuracy by approximately 10% when data availability and label budget are extremely limited. Inner-loop augmentation works well in the extreme case. Nevertheless, the combination of inner- and outer-loop augmentations still improves accuracy.
TABLE IV: Ablation study with different textual augmentation modules of Meta-CAL
@@ VII-C Effect of Augmentations
VII-D Annotation Budgets
@@
VII-E Memory Budgets
We evaluate memory efficiency of our model. Table V compares the performance of three models with the best average accuracy, i.e., our model with Reservoir sampling, our model with Ring Buffer and C-MAML. While all these models use the MAML framework, only the Meta-CAL models employ data augmentation to enhance generalization. The result shows Meta-CAL models outperform C-MAML in all three cases. Notably, Meta-CAL w/. Reservoir can attain more than 50% accuracy by saving only one sample per seen class. As a result, it indicates that data augmentation improves performance.
TABLE V: Accuracy on different memory budgets. The AL strategy is RAND.
@@ VII-E Memory Budgets
Specifically, our model dynamically selects worthwhile unlabeled data to annotate, thereby preventing catastrophic forgetting and maximizing model performance.
VIII Conclusion
This paper considers a realistic continual learning scenario with limited annotation budgets, namely continual active learning. We devise a novel method to address this problem. Our method dynamically selects worthwhile unlabeled data for annotation to maximize model performance and address continual learning problems. Specifically, it applies MAML with memory replay and consistency regularization to enable efficient learning with limited memory and annotation budgets while mitigating catastrophic forgetting. In experiments, we analyzed the effect of various active learning strategies and memory sample selection methods on model performance. The experimental results shows our method achieves a great performance in a 5-shot continual active learning setup. Furthermore, we demonstrate that random sampling is still the best default strategy for both active learning and memory sample selection to solve 5-shot continual active learning problems. However, we only evaluate our method in a 5-shot continual active learning scenario. In the future work, we can extend our evaluation to a more realistic scenario where the amount of labeled data is {100, 1000, 5000}. In addition, the annotation budget is allocated equally for each task. The annotation budget allocation strategies can be further studied.

### v4

@@ 6.5 Main Results
   w/. Ring buffer (5 per class) $\mathbf{66.9}$
@@
   w/. Reservoir [^2] Default memory selection scheme 66.5
@@
FULL [^3] Upper bound performance 64.1 $\pm$ 1.9
As for the memory sample selection schemes, Ring Buffer outperforms Reservoir Sampling by less than 1% in RAND and REP. Whereas, Reservoir Sampling exhibits smaller standard deviations, indicating its robustness to training set orders. Prototype Sampling has the worst performance. We perform further analysis in §7.2.
Table 3: Comparison of Different AL Strategies
@@
44footnotetext: We report the average accuracy of training sequence Yelp $\rightarrow$ AGNews $\rightarrow$ DBpedia $\rightarrow$ Amazon $\rightarrow$ Yahoo.
Table 2 and Table 3 demonstrate that Random is the best AL strategy for our model. As shown in Table 3, for both Diversity- and Representative-based AL strategies, KMeans is the most effective approach. Uncertainty-based methods comparatively perform poorly. CL models aim to revisit past samples to enhance generalization. However, replaying and annotating uncertain samples may not sufficiently improve generalization capabilities, which can hinder effective knowledge consolidation and consequently harms model performance.
7 Further Analysis
7.1 Stability & Plasticity
We test the performance of different AL strategies in terms of stability and plasticity.
(a) RAND
(b) REP (KMeans)
(c) DIV (KMeans)
(d) UNC (Least-confidence)
Stability: We employ the BWT metric as an indicator of catastrophic forgetting, evaluating the impact of the updated model parameters on the performance across all previously learned tasks. The negative BWT values indicate forgetting. As shown in Fig. 2, RAND shows the least forgetting, which indicates the best stability on past tasks. However, all methods exhibit a large forgetting after sequentially learning Task 3. Task 3 (DBpedia, 14 classes) has a relatively large label space compared to Task 1 (Yelp, 5 classes) and Task 2 (AGNews, 4 classes). As shown in Fig. 1, the accuracy on DBpedia exceeds the accuracy on other tasks. Since MAML learns high-quality reusable features for fast adaptation [30], it tends to find shared representations that specifically benefit tasks with large label spaces.
Plasticity: A positive FWT score indicates the model’s ability to leverage knowledge from previously learned tasks, facilitating zero-shot learning and efficient adaptation to the new task. As shown in Fig. 2, RAND, REP, and DIV show positive FWT values after training Task 4. This observation indicates that successful forward knowledge transfer from the previously learned tasks occurs. Fig. 1 further supports that this transfer mainly occurs within the same domain, demonstrating effective domain-incremental learning capabilities. Specifically, the proposed method leverages prior knowledge acquired from a familiar domain to facilitate efficient adaptation to new tasks within that domain.
7.2 Memory Insight
(a) After Learning Task 1
(b) After Learning Task 2
(c) After Learning Task 3
(d) After Learning Task 4
(e) After Learning Task 5
Different learning stages: Fig. 3 presents the T-SNE visualization of memory samples at different learning stages. As show in Figure 3a to 3b, when the proposed model learns a small number of tasks or encounters small label spaces, it focuses on ensuring intra-task generalization. However, as the number of seen tasks increases, the model shifts its focus towards ensuring inter-task generalization as shown in 3d to 3e. Consequently, we observe a phenomenon wherein the memory samples from the last tasks cluster together.
@@ 7.2 Memory Insight
AL strategies: As shown in Fig. 4a to 4d, memory data in RAND show a good data dispersion within the last task (i.e., intra-task generalization) and across multiple tasks (i.e., inter-task generalization), resulting in the best accuracy. In contrast, UNC in Fig. 4d shows subpar performance in inter-task generalization. The result validates UNC’s incapability to achieve generalization and preserve knowledge. Both REP and DIV exhibit lower accuracy compared to RAND, with DIV showing an obvious lower sparsity. Therefore, it is important to find an optimal balance between representativeness and diversity to ensure generalization.
Memory sample selection methods: In Fig. 4a, 4e and 4f, the choice of memory sample selection method also affects the model’s performance. Prototype Sampling selects representative memory samples, potentially resulting in relatively less sparsity across multiple tasks. In contrast, Reservoir and Ring Buffer sampling strategies introduce randomness into the memory sample selection process, achieving better inter-task generalization.
@@
7.3 Effect of Augmentations
We conduct an ablation study to analyze textual augmentations and their key roles in Meta-CAL. We examine the effect of inner-loop augmentation and outer loop augmentation in Table 4. The meta samples in the outer-loop constrain the model behaviour.
Table 4: Ablation Study with Different Augmentation Modules
@@ 7.3 Effect of Augmentations
We hypothesize that if the meta-samples exhibit sufficient generalization capabilities, they can facilitate effective knowledge retention from prior tasks, consequently mitigating the issue of catastrophic forgetting. It is noteworthy that the meta-samples acquired through the combined effect of active learning acquisition and memory sample selection. We further enhance their generalization capabilities through data augmentation. The results validate this hypothesis, as we observe a significant gain by using outer-loop augmentation to enhance generalization. It improves accuracy by approximately 10% when data availability and label budget are extremely limited.
While inner-loop augmentation proves effective in extreme cases, it might not be as advantageous as outer-loop augmentation in non-extreme scenarios. Nevertheless, the combination of inner- and outer-loop augmentations still significantly improves accuracy.
7.4 Annotation Budgets
@@
7.5 Memory Budgets
We evaluate memory efficiency of our model. Table 5 compares the performance of three models with the best average accuracy, i.e., our model with Reservoir sampling, our model with Ring Buffer and C-MAML. While all these models use the MAML framework, only the Meta-CAL models employ consistency regularization to enhance generalization. The result shows Meta-CAL models outperform C-MAML in all three cases. Notably, Meta-CAL w/. Reservoir can attain more than 50% accuracy by saving only one sample per seen class. As a result, it also indicates that consistency regularization through data augmentation improves performance.
Table 5: Accuracy on different memory budgets. The AL strategy is RAND.
@@ 7.5 Memory Budgets
8 Conclusion
This paper considers a realistic continual learning scenario, namely few-shot continual active learning. To address this resource-constrained continual learning problems, we propose a novel method that employing meta-learning and active learning techniques, called Meta-Continual Active Learning. Specifically, our model dynamically queries worthwhile unlabeled data for annotation and reformulate meta-objective with experience rehearsal and consistency regularization. In such a way, the proposed method is able to prevent catastrophic forgetting and improve generalization capability in a low-resource scenario. We conduct extensive experiments on benchmark text classification datasets in a 5-shot continual active learning setting. The results show the robustness of the proposed method. However, we only evaluate our method in a 5-shot case. In the future work, we can extend our evaluation to a more realistic scenario where the amount of labeled data is {100, 1000, 5000} and extend to other NLP tasks, e.g., language model training, text generation, and knowledge base enrichment. Furthermore, the annotation budget is allocated equally for each task. The annotation budget allocation strategies can be further studied.
Appendix A Supplemental Material
Table 6: Accuracy of Four Training Set Orders and Standard Deviation
AL CL Method Order 1 Order 2 Order 3 Order 4 Average
RAND MAML-SEQ $4.1\pm 0.3$ $10.1\pm 0.2$ $14.8\pm 3.6$ $4.4\pm 3.9$ $8.4\pm 5.1$
OML-ER $60.2\pm 3.0$ $59.7\pm 2.1$ $56.4\pm 2.7$ $60.4\pm 4.0$ $59.2\pm 1.9$
C-MAML $61.2\pm 2.1$ $65.1\pm 2.8$ $65.8\pm 1.0$ $59.8\pm 2.3$ $63.0\pm 2.9$
Ours.(PROT) $54.7\pm 1.8$ $61.8\pm 1.6$ $62.1\pm 1.8$ $52.4\pm 2.7$ $57.6\pm 4.9$
Ours.(RB) $66.4\pm 0.5$ $68.1\pm 0.6$ $69.2\pm 0.2$ $63.8\pm 0.2$ $66.9\pm 2.4$
Ours.(RESV) $65.7\pm 0.3$ $68.1\pm 1.2$ $68.1\pm 1.2$ $64.2\pm 0.6$ $66.5\pm 1.9$
MAML-SEQ $10.3\pm 7.1$ $12.2\pm 1.0$ $6.6\pm 2.7$ $1.4\pm 0.1$ $7.63\pm 4.8$
OML-ER $55.5\pm 3.1$ $56.4\pm 1.2$ $58.2\pm 3.4$ $59.3\pm 3.2$ $57.4\pm 1.7$
REP C-MAML $58.3\pm 0.7$ $62.0\pm 2.6$ $64.4\pm 1.8$ $58.9\pm 3.1$ $60.9\pm 2.8$
(KMeans) Ours.(PROT) $58.0\pm 2.8$ $62.3\pm 1.3$ $59.2\pm 3.3$ $51.8\pm 0.6$ $57.8\pm 4.4$
Ours.(RB) $65.3\pm 1.0$ $67.3\pm 0.6$ $68.7\pm 1.4$ $63.7\pm 1.5$ $66.3\pm 2.2$
Ours.(RESV) $65.2\pm 0.5$ $67.5\pm 0.9$ $67.3\pm 0.2$ $63.5\pm 1.2$ $65.9\pm 1.9$
MAML-SEQ $7.3\pm 2.2$ $9.6\pm 1.6$ $10.5\pm 6.3$ $1.5\pm 0.1$ $7.23\pm 4.1$
OML-ER $54.8\pm 1.8$ $54.5\pm 1.2$ $59.1\pm 2.8$ $60.8\pm 3.1$ $57.3\pm 3.1$
DIV C-MAML $60.9\pm 1.7$ $59.1\pm 2.1$ $60.6\pm 1.7$ $59.1\pm 2.1$ $59.9\pm 1.0$
(KMeans) Ours.(PROT) $55.5\pm 3.2$ $63.4\pm 1.4$ $59.7\pm 1.1$ $52.6\pm 4.6$ $57.8\pm 4.7$
Ours.(RB) $63.7\pm 0.4$ $66.8\pm 1.5$ $66.8\pm 0.3$ $61.7\pm 2.1$ $64.8\pm 2.5$
Ours.(RESV) $64.7\pm 1.1$ $65.7\pm 0.6$ $66.1\pm 0.2$ $63.1\pm 2.3$ $64.9\pm 1.3$
MAML-SEQ $4.2\pm 2.0$ $9.2\pm 1.8$ $9.6\pm 4.3$ $1.6\pm 0.4$ $6.15\pm 3.9$
OML-ER $47.1\pm 1.5$ $51.9\pm 2.8$ $46.7\pm 7.6$ $44.0\pm 1.5$ $47.4\pm 3.3$
UNC C-MAML $50.9\pm 3.3$ $49.2\pm 1.5$ $53.1\pm 2.5$ $48.4\pm 0.4$ $50.4\pm 2.1$
(L-CONF) Ours.(PROT) $37.5\pm 6.5$ $49.5\pm 0.9$ $44.6\pm 6.0$ $36.3\pm 3.5$ $42.0\pm 6.2$
Ours.(RB) $47.4\pm 3.6$ $53.9\pm 1.1$ $53.7\pm 2.0$ $47.8\pm 1.4$ $50.7\pm 3.6$
Ours.(RESV) $53.3\pm 3.8$ $54.4\pm 0.5$ $56.3\pm 1.3$ $49.8\pm 0.8$ $53.5\pm 2.7$
REP(Means) Ours.(RESV) $57.5\pm 0.9$ $58.1\pm 0.4$ $60.6\pm 1.2$ $55.5\pm 1.0$ $57.9\pm 2.1$
DIV(Means) $50.0\pm 1.5$ $50.5\pm 2.8$ $52.5\pm 0.4$ $49.6\pm 4.7$ $50.7\pm 1.3$
UNC(MARG) $58.4\pm 1.8$ $58.0\pm 0.9$ $59.1\pm 2.1$ $49.2\pm 1.4$ $56.2\pm 4.7$
UNC(ENT) $51.2\pm 0.3$ $49.2\pm 1.4$ $54.6\pm 2.3$ $48.5\pm 4.1$ $50.9\pm 2.7$
FULL C-MAML $64.5\pm 1.4$ $62.2\pm 1.3$ $66.5\pm 1.0$ $63.2\pm 1.7$ $64.1\pm 1.7$